# SlipStream: Adapting Pipelines for Distributed Training of Large DNNs Amid Failures

### v1

@@ 6. Related Work
Parallel Training. Data parallelism (Krizhevsky et al., 2017; Chen et al., 2016; Dean et al., 2012) is a widely utilized mode of parallelism that distributes the dataset across different partitions for processing. In this mode, the learned weights are synchronized via either an all-reduce approach (Chen et al., 2016) or by using parameter servers (Li et al., 2014; Chilimbi et al., 2014; Jiang et al., 2020). Alternatively, model parallelism (Krizhevsky, 2014; Shoeybi et al., 2019; Shazeer et al., 2018) involves distributing the components of a deep neural network (DNN) model across multiple GPU devices, allowing each device to handle a specific subset of the model’s parameters for all input data. Recently, pipeline parallelism (Huang et al., 2019; Narayanan et al., 2019; Li and Hoefler, 2021; Liu et al., 2023; Kim et al., 2023) has emerged as a technique for training large models by dividing the model’s layers among different workers and utilizing microbatches to efficiently use the available computational resources. Prominent deep learning training frameworks like PyTorch (PyTorch, 2024), DeepSpeed (Rasley et al., 2020), and Megatron (Narayanan et al., 2021) have adopted hybrid-parallelism, an approach that integrates data parallelism, model parallelism, and pipeline parallelism. This integration facilitates training on a massive scale while enhancing computational and memory efficiency. Additionally, DeepSpeed introduces ZeRO-style (Rajbhandari et al., 2020, 2021; Ren et al., 2021; Zhao et al., 2023; Korthikanti et al., 2023) data parallelism, which strategically partitions model states across GPUs, coordinating via communication collectives to synchronize parameters as needed.
@@
Optimizing Hybrid-Parallel Training. Improvements in hybrid parallelism are extensively studied in the context of ML training (Narayanan et al., 2021; Korthikanti et al., 2023; Rasley et al., 2020; Li et al., 2020; Zhao et al., 2023; Zheng et al., 2022; Huang et al., 2019; Narayanan et al., 2019; Li and Hoefler, 2021; Liu et al., 2023; Kim et al., 2023). Alpa (Zheng et al., 2022) automatically optimizes inter- and intra-operator parallelism using a hierarchical ILP formulation. Tofu (Wang et al., 2019a) uses DP to optimally partition tensor operations in a single node. FlexFlow (Jia et al., 2018) uses a randomized search algorithm to quickly find parallelism strategies. TensorOpt (Cai et al., 2022) introduces a DP that can optimize parallelism for multiple resource dimensions (e.g., memory and compute). Piper (Tarnawski et al., 2021) proposes a two-level DP algorithm to find optimal hybrid parallelism strategies. SlipStream uses a DP algorithm and an MILP which exploit key functional redundancies in hybrid parallelism to explicitly improve training throughput under failures.
@@
Elastic Training. Elastic training systems can dynamically adjust the resources allocated to a training job. Both Horovod (Sergeev and Del Balso, 2018) and Torch Distributed (Li et al., 2020) provide mechanisms to modify the number of workers, but require either a restart from a checkpoint or an expensive re-shuffling of model parameters. CoDDL (Hwang et al., 2021), Optimus (Peng et al., 2018), OASiS (Bao et al., 2018), and Themis (Mahajan et al., 2020) are ML cluster schedulers that can dynamically allocate resources across multiple DNN training jobs. Or et al. (Or et al., 2020) auto-scale the number of workers for a training job. Varuna (Athlur et al., 2022) provides elastic training by leveraging spot instances, but requires restarts from checkpoints to handle unexpected pre-emptions. While some elastic training systems can affect model consistency by changing important hyperparameters such as batch size and learning rate, SlipStream ensures that operations are mathematically consistent regardless of the number of failures.
@@
Fault-tolerant Training. DNN training systems widely leverage checkpoints as the basic mechanism for fault recovery (Rasley et al., 2020; Zu et al., 2024; Le Scao et al., 2022; Jiang et al., 2024; He et al., 2023; Beaumont, 2022; Jeon et al., 2019; Wang et al., 2019b). While naive checkpointing can lead to training stalls, recent work aims to reduce these overheads. CheckFreq (Mohan et al., 2021) dynamically adapts the checkpointing frequency to reduce overheads. Check-N-Run (Eisenman et al., 2022) quantizes in-frequently updated embedding table checkpoints for recommendation models. Gemini (Wang et al., 2023) introduces checkpoint traffic scheduling across the storage hierarchy (local CPU memory, remote CPU memory, and remote persistent storage). Megascale (Jiang et al., 2024) alleviates the storage bottleneck during recovery by sharing data between corresponding GPU workers across data parallel groups.
@@
Two recent projects have proposed to utilize pipeline parallelism for efficient training in the presence of failures without spares. Bamboo (Thorpe et al., 2023) introduces redundant computation (RC) to provide resilience in the presence of frequent preemptions of training with spot instances. Oobleck (Jang et al., 2023) provides resiliency through creation of heterogenous pipelines. SlipStream matches or outperforms them for a wide range of model sizes and failure frequencies.
@@ 7. Conclusion
SlipStream enables efficient DNN training in the presence of failures without the use of spare resources. SlipStream exploits the functional redundancies inherent to hybrid parallel training systems used for large DNNs. It introduces novel techniques to minimize the training throughput degradation of sustaining failures by exploiting unused resources (e.g., pipeline bubbles). We evaluated an end-to-end SlipStream prototype and showed it can tolerate a high numbers of concurrent failures for systems and models of various sizes. SlipStream improves training throughput under failures over Oobleck and Bamboo by up to 1.46×1.46$\times$1.46 × and 1.64×1.64$\times$1.64 × respectively.

### v2

@@ 7. Related Work
Parallel Training. Data parallelism (Krizhevsky et al., 2017; Chen et al., 2016; Dean et al., 2012) is a widely utilized mode of parallelism that distributes the dataset across different partitions for processing. In this mode, the learned weights are synchronized via either an all-reduce approach (Chen et al., 2016) or by using parameter servers (Li et al., 2014; Chilimbi et al., 2014; Jiang et al., 2020). Alternatively, model parallelism (Krizhevsky, 2014; Shoeybi et al., 2019; Shazeer et al., 2018) involves distributing the components of a deep neural network (DNN) model across multiple GPU devices, allowing each device to handle a specific subset of the model’s parameters for all input data. Recently, pipeline parallelism (Huang et al., 2019; Narayanan et al., 2019; Li and Hoefler, 2021; Liu et al., 2023; Kim et al., 2023) has emerged as a technique for training large models by dividing the model’s layers among different workers and utilizing micro-batches to efficiently use the available computational resources. Prominent deep learning training frameworks like PyTorch (pyt, 2024), DeepSpeed (Rasley et al., 2020), and Megatron (Narayanan et al., 2021) have adopted hybrid-parallelism, an approach that integrates data parallelism, model parallelism, and pipeline parallelism. This integration facilitates training on a massive scale while enhancing computational and memory efficiency. Additionally, DeepSpeed introduces ZeRO-style (Rajbhandari et al., 2020, 2021; Ren et al., 2021; Zhao et al., 2023; Korthikanti et al., 2023) data parallelism, which strategically partitions model states across GPUs, coordinating via communication collectives to synchronize parameters as needed.
@@
Optimizing Hybrid-Parallel Training. Improvements in hybrid parallelism are extensively studied in the context of ML training (Narayanan et al., 2021; Korthikanti et al., 2023; Rasley et al., 2020; Li et al., 2020; Zhao et al., 2023; Zheng et al., 2022; Huang et al., 2019; Narayanan et al., 2019; Li and Hoefler, 2021; Liu et al., 2023; Kim et al., 2023). Alpa (Zheng et al., 2022) automatically optimizes inter- and intra-operator parallelism using a hierarchical ILP formulation. Tofu (Wang et al., 2019a) employs dynamic programming to optimally partition tensor operations in a single node. FlexFlow (Jia et al., 2018) uses a randomized search algorithm to quickly find parallelism strategies. TensorOpt (Cai et al., 2022) introduces a dynamic programming approach capable of optimizing parallelism across multiple resource dimensions, including memory and compute. Piper (Tarnawski et al., 2021) proposes a two-level dynamic programming algorithm to find optimal hybrid parallelism strategies. In contrast, ReCycle leverages a dynamic programming algorithm and mixed-integer linear programming (MILP) to exploit key functional redundancies in hybrid parallelism, explicitly enhancing training throughput in the presence of failures.
@@
Elastic Training. Elastic training systems can dynamically adjust the resources allocated to a training job. Both Horovod (Sergeev and Del Balso, 2018) and Torch Distributed (Li et al., 2020) offer mechanisms to modify the number of workers, but require either a restart from a checkpoint or an expensive re-shuffling of model parameters. CoDDL (Hwang et al., 2021), Optimus (Peng et al., 2018), OASiS (Bao et al., 2018), and Themis (Mahajan et al., 2020) are ML cluster schedulers that can dynamically allocate resources across multiple DNN training jobs. Or et al. (Or et al., 2020) auto-scale the number of workers for a training job. Varuna (Athlur et al., 2022) provides elastic training by leveraging spot instances, but requires restarts from checkpoints to handle unexpected preemptions. While some elastic training systems can affect model consistency by changing critical hyperparameters such as batch size and learning rate, ReCycle guarantees mathematical consistency of operations, regardless of the number of failures.
@@
Fault-tolerant Training. DNN training systems commonly use checkpoints for fault recovery (Rasley et al., 2020; Zu et al., 2024; Le Scao et al., 2022; Jiang et al., 2024; He et al., 2023; Beaumont, 2022; Jeon et al., 2019; Wang et al., 2019b), though naive checkpointing can cause stalls. Recent works like CheckFreq (Mohan et al., 2021), Check-N-Run (Eisenman et al., 2022), and Gemini (Wang et al., 2023) reduce overheads by adapting checkpoint frequency, quantizing embedding tables, and scheduling checkpoint traffic across the storage hierarchy, respectively. Megascale (Jiang et al., 2024) alleviates the storage bottleneck during recovery by sharing data between corresponding GPU workers across data parallel groups.
@@
Two recent projects have proposed to utilize pipeline parallelism for efficient training in the presence of failures without spares. Bamboo (Thorpe et al., 2023) introduces redundant computation (RC) to provide resilience in the presence of frequent preemptions of training with spot instances. Oobleck (Jang et al., 2023) provides resiliency through creation of heterogeneous pipelines. ReCycle matches or outperforms them for a wide range of model sizes and failure frequencies.
@@ 8. Conclusion
ReCycle enables efficient DNN training in the presence of failures without relying on spare resources. By leveraging the functional redundancies inherent in hybrid parallel training systems for large DNNs, ReCycle introduces innovative techniques to minimize the degradation in training throughput caused by failures. It does this by utilizing unused resources, such as pipeline bubbles. We evaluated an end-to-end prototype of ReCycle and demonstrated its ability to tolerate a high number of concurrent failures across systems and models of varying sizes. Compared to Oobleck and Bamboo, ReCycle improves training throughput under failure conditions by up to 1.46×1.46$\times$1.46 × and 1.64×1.64$\times$1.64 × respectively.
Acknowledgements
We are grateful to the anonymous reviewers and to our shepherd, Laurent Bindschaedler, whose comments have greatly helped improve this paper. This research was partly supported by the Stanford Platform Lab and its affiliates, and by ACE, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA. Athinagoras Skiadopoulos was supported by a Stanford Graduate Fellowship. Mark Zhao was supported by a Stanford Graduate Fellowship and a Meta PhD Fellowship.